# Localized Multi-Dimensional Patterns

### v1

@@ 6.1 Localized (1+2)D Patterns
where $a_{1},a_{2}$ are fixed constants, while the ‘nonlocal’ coefficients $\mu,\gamma,b$ depend on the spatial average of $A$. Furthermore, localized solutions no longer require rolls to bifurcate subcritically in order to emerge and persist outside the bistability region. Localized $(1+2)$D patterns have also been observed in turbulence of plane Couette flow [211, 212, 193]; however, these solutions have a strong connection to chaos, and so they are extremely difficult to study using deterministic methods [13].
@@ 6.2 Localized (2+1)D patterns
Localized planar patterns have been observed in various hydrodynamic settings, including binary fluid convection [162], magnetohydrodynamics [35, 120], and turbulence of plane Couette flow [46]. Between these examples, the only thorough investigation of the bifurcation structure of localized planar patterns comes from the work of Lo Jacono et al. [161, 162] on binary fluid convection. Here, the authors use spectral element methods to numerically identify localized rhombic, i.e. $\mathbb{D}_{2}$-symmetric, and square, i.e. $\mathbb{D}_{4}$-symmetric, patterns that are then continued in a system parameter to trace out their bifurcation diagrams. Localized planar patterns with both $\mathbb{D}_{2}$ and $\mathbb{D}_{4}$ symmetries are found to emerge and undergo snaking, where the pattern grows in width. However, in contrast to the localized dihedral patterns observed in [115, 155], the patterns do not grow in a regular fashion by maintaining a convex core of periodic pattern. Rather, as one can see in Figure 47, grow takes place through the extension of rolls in a symmetric arm-like configuration. The ‘arms’ of each localized planar pattern exhibit similar behavior to the worm solutions studied by Avitabile et al. [16] for the cubic-quintic SHE. Lo Jacono et al. further investigated localized planar patterns in binary fluid convection in various sizes of domains [162]. They observed that, for relatively small domains, localized rhombic and square patterns grow until they fill the domain, generating spatially extended convection patterns. In contrast, if the domain is large enough, each pattern maintains a ‘convex core - extended arms’ structure as observed previously. As such, the pattern never fills the domain and does not generate any kind of spatially extended convection pattern, as shown in Figure 47.
@@
for a complex field $\zeta=\zeta(x,y)\in\mathbb{C}$. Localized planar patterns are then approximated by a unique ground state of (6.2). Figure 48 provides an illustration of such minimizers. For a more detailed review of the functional analytic approaches for localized planar patterns in water waves see [112, Section 8.3].
@@ 7 Conclusion and Open Problems
The study of localized patterns has seen significant development over the past 30 years. It has motivated new mathematical methods and techniques that can be used for a variety of applications that go far beyond localized patterns. The use of numerical continuation and the increase in speed of modern computers has allowed researchers to uncover new phenomena and provide evidence for interesting mathematical questions. It is clear the understanding of one-dimensional localized patterns is strongest, while the community’s grasp weakens as the dimension of the patterns arise on increases.
@@
While it is clear from the review that developing mathematical theory for 3D localized patterns is a major and fruitful area for future research, there exists other avenues that might be more tractable in the immediate future. Many have been highlighted as we progressed through the individual sections, but here we enumerate some of the largest outstanding problems that would significantly advance the field.
@@
    Stability analysis of the 3D localized patterns in §6 is very much an open problem. This could potentially be tackled through a combination of variational and spatial dynamical methods. Spatial inhomogeneities can have a significant impact on pattern selection, and there are some recent advances in this direction [129, 130].

### v2

@@ 6.1 Localized (1+2)D Patterns
where $a_{1},a_{2}$ are fixed constants, while the ‘nonlocal’ coefficients $\mu,\gamma,b$ depend on the spatial average of $A$. Furthermore, localized solutions no longer require SPPs to bifurcate subcritically in order to emerge and persist outside the bistability region. Localized $(1+2)$D patterns have also been observed in turbulence of plane Couette flow [218, 219, 200]; however, these solutions have a strong connection to chaos, and so they are extremely difficult to study using deterministic methods [13].
@@ 6.2 Localized (2+1)D patterns
Localized planar patterns have been observed in various hydrodynamic settings, including binary fluid convection [167], magnetohydrodynamics [35, 124], and turbulence of plane Couette flow [46]. Between these examples, the only thorough investigation of the bifurcation structure of localized planar patterns comes from the work of Lo Jacono et al. [166, 167] on binary fluid convection. Here, the authors use spectral element methods to numerically identify localized rhombic, i.e. $\mathbb{D}_{2}$-symmetric, and square, i.e. $\mathbb{D}_{4}$-symmetric, patterns that are then continued in a system parameter to trace out their bifurcation diagrams. Localized planar patterns with both $\mathbb{D}_{2}$ and $\mathbb{D}_{4}$ symmetries are found to emerge and undergo snaking, where the pattern grows in width. However, in contrast to the localized dihedral patterns observed in [119, 160], the patterns do not grow in a regular fashion by maintaining a convex core of periodic pattern. Rather, as one can see in Figure 47, grow takes place through the extension of SPPs in a symmetric arm-like configuration. The ‘arms’ of each localized planar pattern exhibit similar behavior to the worm solutions studied by Avitabile et al. [16] for the cubic-quintic SHE. Lo Jacono et al. further investigated localized planar patterns in binary fluid convection in various sizes of domains [167]. They observed that, for relatively small domains, localized rhombic and square patterns grow until they fill the domain, generating spatially extended convection patterns. In contrast, if the domain is large enough, each pattern maintains a ‘convex core - extended arms’ structure as observed previously. As such, the pattern never fills the domain and does not generate any kind of spatially extended convection pattern, as shown in Figure 47.
@@
for a complex field $\zeta=\zeta(x,y)\in\mathbb{C}$. Localized planar patterns are then approximated by a unique ground state of (6.2). Figure 48 provides an illustration of such minimizers. For a more detailed review of the functional analytic approaches for localized planar patterns in water waves see [116, Section 8.3].
@@ 7 Conclusion and Open Problems
The study of localized patterns has seen significant development over the past 30 years. It has motivated new mathematical methods and techniques that can be used for a variety of applications that go far beyond localized patterns. The use of numerical continuation and the increase in speed of modern computers has allowed researchers to uncover new phenomena and provide evidence for interesting mathematical questions. It is clear the understanding of one-dimensional localized patterns is strongest and most complete, while the community’s grasp weakens as the dimension the patterns arise on increases.
@@
While it is clear from the review that developing mathematical theory for 3D localized patterns is a major and fruitful area for future research, there exists other avenues that might be more tractable in the immediate future. Many have been highlighted as we progressed through the individual sections, but here we enumerate some outstanding problems that would significantly advance the field.
@@
    Developing a general stability theory in particular for 3D localized patterns in §6 is very much an open problem. This could potentially be tackled through a combination of variational and spatial dynamical methods. Spatial inhomogeneities can have a significant impact on pattern selection, and there are some recent advances in this direction [134, 135].